# Transverse Momentum-Dependent Heavy-Quark Fragmentation at Next-to-Leading Order

### v1

@@ 7.1 Mass-dependent EEC jet functions
In this section we use our results to produce explicit results for the the effect of heavy-quarks on the energy-energy correlator (EEC) observable [70] as measured in $e^{+}e^{-}$ collisions,
@@ 7.4 Extension to nonperturbative $1-z_{\chi,H}\sim\Lambda_{\mathrm{QCD}}^{2}/Q^{2}$
Recently, an independent NLO calculation of the heavy-quark TMD FFs $d_{1\,Q/Q}$ and $d_{1\,Q/g}$ was performed in ref. [13]. Before commenting on the differences between the two calculational approaches, we first want to stress that after accounting for a measure factor of $1/z^{2}$ and the renormalized TMD soft function in eq. (B.4), we find complete agreement of our eqs. (3.4) and (3.5) with the final $b_{T}$-space results given in their eqs. (2.36) and (4.26). In order to gauge the strength of this cross check, it is relevant to note that while the calculation of the real-emission diagrams in ref. [13], like ours in section 3.1, was initially performed in momentum space and only Fourier-transformed after renormalization, ref. [13] used a completely orthogonal method to isolate the singularities in $d^{(1)}_{1\,Q/Q}$, i.e., to address the key source of subtleties. Specifically, in ref. [13] the characteristic denominator $1/X\equiv 1/(k_{T}^{2}+m^{2}\bar{z}^{2}/z^{2})$ appearing in the real-emission diagrams was directly expanded into one-dimensional plus distribution of $X$ and in addition was multiplied by one-dimensional distributions of $1-z$, which effectively amounts to performing a change of variables to $(z,X)$. This procedure is completely orthogonal to our approach, namely, we expand in terms of one-dimensional plus distributions along the boundaries $\bar{z}=0$ and $k_{T}=0$ and a two-dimensional plus distribution whose defining domain aligns with the main variables $(z,k_{T})$. The two approaches in particular differ (and have to differ) for the boundary contribution proportional to $\delta(1-z)$. (Indeed, the process in ref. [13] generates highly nontrivial twofold integrals $F(m^{2}/\Lambda^{2})$ and $G(m^{2}/\Lambda^{2})$ multiplying $\delta(1-z)$, where $\Lambda$ is the boundary condition of the $X$ plus distributions, with no closed form for $F$ and $G$ away from asymptotic limits for $m$.)
@@
One difference between the two power-counting assumptions on $1-z_{H}$ regards the kind of cross check the respective bHQET calculation provides on the large-mass behavior of the TMD FF. The approach of ref. [13] checks the differential $z$ dependence of $d_{1\,Q/Q}^{(1)}$, but in intermediate steps an identical plus distribution expansion with respect to a very similar, complicated denominator $1/X^{\prime}=1/(k_{T}^{2}+\bar{z}^{2}m^{2})$, which is subject to the same subtleties as the full calculation. In this sense, our calculation provides a strong complementary check of the large-mass behavior since it directly probes the total integral, i.e., a defining property of the distributional result. [^5] Ref. [13] also performed a small-mass consistency check, using the one-loop TMD PDF matching coefficients and a reciprocity relation. This provides a sufficient check at one loop, but reciprocity relations between TMD PDF and FF matching coefficients are known to be subtle at higher orders [47, 46]. Here we stress that the correct all-order relation for the small-mass limit is the one we originally gave in ref. [1], i.e., it has to involve the genuine TMD FF matching coefficients onto twist-2 collinear FFs because the UV matching coefficients are independent of the details of the IR, that is, whether one matches onto a theory at $\mu\sim\Lambda_{\mathrm{QCD}}$ or $\mu\sim m$.
A second and final difference between the power counting assumptions is phenomenological and regards the kinds of differential distributions one may describe in either approach. Our stance here is that while – again – the assumption of ref. [13] is sufficient and entirely correct to describe specific regions of the multi-differential $(z_{H},k_{T})$ distribution, it is not the unique power-counting assumption on $1-z_{H}$, $k_{T}$, $m$ that would yield a leading-power factorization theorem involving bHQET modes, in particular not when different possible scalings of $\Lambda_{\mathrm{QCD}}$ are also considered. As one example, we point to our analysis in section 6, where we continued to count $m\sim k_{T}$, but still took $1-z_{H}$ small, and likewise obtained a leading-power factorization theorem, in that case also involving a purely collinear bHQET shape function, but nontrivial $k_{T}$-dependent correlators of lightlike Wilson lines at the scale $\mu\sim k_{T}$. In light of this, we believe that capturing the full $(z_{H},k_{T})$ dependence will require dedicated future studies combining all of these results, while for the time being and within the scope of exploratory theory studies (or, necessarily less differential, initial experimental measurements) the restriction to wide bins or total integrals over $z_{H}$ is reasonable.
@@ Acknowledgments
We thank Iain Stewart and Kyle Lee for helpful discussions. We thank the Erwin Schrödinger Institute for hospitality while parts of this work were performed. This work was supported in part by the Office of Nuclear Physics of the U.S. Department of Energy under Contract No. DE-SC0011090. Z.S. was also supported by a fellowship from the MIT Department of Physics. R.v.K. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002090 COLORFREE). J.M. was supported by the D-ITP consortium, a program of NWO that is funded by the Dutch Ministry of Education, Culture and Science (OCW).
@@ Appendix B Perturbative ingredients
The combined collinear and soft mass matching coefficient $C_{m}$, see ref. [1] for our notation, is given by [43]
@@
The $\zeta$ dependence arises from the rapidity renormalization of secondary soft heavy-quark contributions starting at $\mathcal{O}(\alpha_{s}^{2})$. The perturbative HQET shape function is given by [5, 6]
@@
$\sqrt{S}(b_{T},\mu,\nu)=1+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-L_{b}^{2}+4L_{b}\ln\frac{\mu}{\nu}-\frac{\pi^{2}}{6}\Bigr{]}\,.$ (B.4)
@@
$\mathcal{S}(\ell^{-},b_{T},\mu,\nu)=\delta(\ell^{-})+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-8L_{b}\,\mathcal{L}_{0}(\ell^{-},\nu)\Bigr{]}\,.$ (B.5)
@@ Appendix C Results for Mellin moments
$\int_{0}^{1}\!\mathrm{d}z\,z^{N}\,\bigl{[}f(x,z)\bigr{]}_{+,+}=\Bigl{[}\int_{0}^{1}\!\mathrm{d}z\,\bigl{(}z^{N}-1\bigr{)}f(x,z)\Bigr{]}\,,$ (C.2)
@@
Refer to captionRefer to caption
Figure -1074: Left: Numerical Mellin moments $\mathcal{R}^{[N]}(y)$ of the finite remainder term in $d_{1\,Q/Q}$. Right: Numerical Mellin moments $d_{1\,Q/g}^{[N]\,(1)}$ of the gluon-to-massive-quark TMD FF.
@@
where $\psi(x)$ ($\psi^{(1)}(x)$) is (the first derivative of) the Digamma function. We were not able to find a closed-form expression for the Mellin transform ~$\tilde{\mathcal{R}}^{[N]}(y)\equiv\int_{0}^{1}\!\mathrm{d}z\,z^{N}\,\tilde{\mathcal{R}}(z,y)$ of the finite remainder term, but implementing the $z$ integral of eq. (3.4) directly leads to stable numerical results (shown in the left panel of figure -1074 for reference) that could readily be interpolated. For $\mathcal{J}_{g/Q}$, the Mellin transform for general $N$ is trivial since $z$ does not appear in the arguments of the Bessel $K$ functions,
@@
For $d_{1\,Q/g}^{[N]\,(1)}$, a closed-form expression in terms of ${}_{1}F_{2}$ hypergeometric and Meijer $G$ functions may be obtained for general $N$ from its definition, but is no more instructive than numerical $z$ moments of eq. (3.5). We show the latter in the right panel of figure -1074 for reference.

### v3

@@ 7.1 Mass-dependent EEC jet functions
In this section we use our results to produce explicit results for the effect of heavy quarks on the energy-energy correlator (EEC) observable [76] as measured in $e^{+}e^{-}$ collisions,
@@ 7.4 Extension to nonperturbative $1-z_{\chi,H}\sim\Lambda_{\mathrm{QCD}}^{2}/Q^{2}$
Recently, an independent NLO calculation of the heavy-quark TMD FFs $d_{1\,Q/Q}$ and $d_{1\,Q/g}$ was performed in ref. [13]. Before commenting on the differences between the two calculational approaches, we first want to stress that after accounting for a measure factor of $1/z^{2}$ and the renormalized TMD soft function in eq. (B.4), we find complete agreement of our eqs. (3.4) and (3.5) with the final $b_{T}$-space results given in their eqs. (2.36) and (4.26). In order to gauge the strength of this cross check, it is relevant to note that while the calculation of the real-emission diagrams in ref. [13], like ours in section 3.1, was initially performed in momentum space and only Fourier-transformed after renormalization, ref. [13] used a completely orthogonal method to isolate the singularities in $d^{(1)}_{1\,Q/Q}$, i.e., to address the key source of subtleties. Specifically, in ref. [13] the characteristic denominator $1/X\equiv 1/(k_{T}^{2}+m^{2}\bar{z}^{2}/z^{2})$ appearing in the real-emission diagrams was directly expanded into one-dimensional plus distributions of $X$ and in addition was multiplied by one-dimensional distributions of $1-z$, which effectively amounts to performing a change of variables to $(z,X)$. This procedure is completely orthogonal to our approach, namely, we expand in terms of one-dimensional plus distributions along the boundaries $\bar{z}=0$ and $k_{T}=0$ and a two-dimensional plus distribution whose defining domain aligns with the main variables $(z,k_{T})$. The two approaches in particular differ (and have to differ) for the boundary contribution proportional to $\delta(1-z)$. Indeed, the process in ref. [13] generates highly nontrivial twofold integrals $F(m^{2}/\Lambda^{2})$ and $G(m^{2}/\Lambda^{2})$ multiplying $\delta(1-z)$, where $\Lambda$ is the boundary condition of the $X$ plus distributions, with no closed form for $F$ and $G$ away from asymptotic limits in $m$.
@@
One difference between the two power-counting assumptions on $1-z_{H}$ regards the kind of cross check the respective bHQET calculation provides on the large-mass behavior of the TMD FF. The approach of ref. [13] checks the differential $z$ dependence of $d_{1\,Q/Q}^{(1)}$, but in intermediate steps an identical plus distribution expansion with respect to a very similar, complicated denominator $1/X^{\prime}=1/(k_{T}^{2}+\bar{z}^{2}m^{2})$ is used, which is subject to the same subtleties as the full calculation. In this sense, our calculation provides a strong complementary check of the large-mass behavior since it directly probes the total integral, i.e., a defining property of the distributional result. [^6] Ref. [13] also performed a small-mass consistency check, using the one-loop TMD PDF matching coefficients and a reciprocity relation. This provides a sufficient check at one loop, but reciprocity relations between TMD PDF and FF matching coefficients are known to be subtle at higher orders [51, 50]. Here we stress that the correct all-order relation for the small-mass limit is the one we originally gave in ref. [1], i.e., it has to involve the genuine TMD FF matching coefficients onto twist-2 collinear FFs because the UV matching coefficients are independent of the details of the IR, that is, whether one matches onto a theory at $\mu\sim\Lambda_{\mathrm{QCD}}$ or $\mu\sim m$.
A second and final difference between the power counting assumptions is phenomenological and regards the kinds of differential distributions one may describe in either approach. Our stance here is that while – again – the assumption of ref. [13] is sufficient and entirely correct to describe specific regions of the multi-differential $(z_{H},k_{T})$ distribution, it is not the unique power-counting assumption on $1-z_{H}$, $k_{T}$, and $m$ that would yield a leading-power factorization theorem involving bHQET modes, in particular not when different possible scalings of $\Lambda_{\mathrm{QCD}}$ are also considered. As one example, we point to our analysis in section 6, where we continued to count $m\sim k_{T}$, but still took $1-z_{H}$ small, and likewise obtained a leading-power factorization theorem, in that case also involving a purely collinear bHQET shape function, but nontrivial $k_{T}$-dependent correlators of lightlike Wilson lines at the scale $\mu\sim k_{T}$. In light of this, we believe that capturing the full $(z_{H},k_{T})$ dependence will require dedicated future studies combining all of these results, while for the time being and within the scope of exploratory theory studies (or, necessarily less differential, initial experimental measurements) the restriction to wide bins or low-moment integrals over $z_{H}$ is reasonable.
@@ Acknowledgments
We thank Iain Stewart, Kyle Lee, and Maximilian Stahlhofen for helpful discussions. We thank the Erwin Schrödinger Institute for hospitality while parts of this work were performed. This work was supported in part by the Office of Nuclear Physics of the U.S. Department of Energy under Contract No. DE-SC0011090. Z.S. was also supported by a fellowship from the MIT Department of Physics. R.v.K. was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101002090 COLORFREE). J.M. was supported by the D-ITP consortium, a program of NWO that is funded by the Dutch Ministry of Education, Culture and Science (OCW).
@@ Appendix B Perturbative ingredients
The combined collinear and soft mass matching coefficient $C_{m}$ see also ref. [1] for our notation, is given by [47]
@@
At this order, it is also equal to the matching coefficient $H_{m,n}^{Q\to h_{v}}(m,\mu,\nu/\sqrt{\zeta})$ (defined in ref. [47], without the superscript) for matching massive collinear quark modes with label momentum $\sqrt{\zeta}$ onto bHQET on their own, whereas the additional contribution $\sqrt{H_{m,s}}(m,\mu,\nu)$ from matching secondary soft heavy quarks onto light soft modes becomes nontrivial at $\mathcal{O}(\alpha_{s}^{2})$. Two-loop expressions for both are given in ref. [47], where they were extracted from the NNLO heavy-quark form factor [79, 80]. At that order, there is also a nonzero $\zeta$ dependence that arises from the rapidity renormalization of $H_{m,n}^{Q\to h_{v}}$ and $H_{m,s}$. The two-loop expression for secondary heavy-quark effects on the light-to-light collinear matching coefficient $H_{m,n}^{q\to q}$ is given in ref. [48] as $H_{c}$, where it was extracted from the results of refs. [81, 47]. The perturbative HQET shape function is given by [5, 6]
@@
$\sqrt{S}(b_{T},m,\mu,\nu)=1+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-L_{b}^{2}+4L_{b}\ln\frac{\mu}{\nu}-\frac{\pi^{2}}{6}\Bigr{]}\,.$ (B.4)
@@
$\mathcal{S}(\ell^{-},b_{T},m,\mu,\nu)=\delta(\ell^{-})+\frac{\alpha_{s}C_{F}}{4\pi}\Bigl{[}-8L_{b}\,\mathcal{L}_{0}(\ell^{-},\nu)\Bigr{]}\,.$ (B.5)
@@ Appendix C Results for Mellin moments
$\int_{0}^{1}\!\mathrm{d}z\,z^{N}\,\bigl{[}f(x,z)\bigr{]}_{+,+}=\Bigl{[}\int_{0}^{1}\!\mathrm{d}z\,\bigl{(}z^{N}-1\bigr{)}f(x,z)\Bigr{]}_{+}\,,$ (C.2)
@@
Figure -1194: Left: Numerical Mellin moments $\mathcal{R}^{[N]}(y)$ of the finite remainder term in $d_{1\,Q/Q}$. Right: Numerical Mellin moments $d_{1\,Q/g}^{[N]\,(1)}$ of the gluon-to-massive-quark TMD FF.
@@
where $\psi(x)$ ($\psi^{(1)}(x)$) is (the first derivative of) the Digamma function. We were not able to find a closed-form expression for the Mellin transform ~$\tilde{\mathcal{R}}^{[N]}(y)\equiv\int_{0}^{1}\!\mathrm{d}z\,z^{N}\,\tilde{\mathcal{R}}(z,y)$ of the finite remainder term, but implementing the $z$ integral of eq. (3.4) directly leads to stable numerical results (shown in the left panel of figure -1194 for reference) that could readily be interpolated. For $\mathcal{J}_{g/Q}$, the Mellin transform for general $N$ is trivial since $z$ does not appear in the arguments of the Bessel $K$ functions,
@@
For $d_{1\,Q/g}^{[N]\,(1)}$, a closed-form expression in terms of ${}_{1}F_{2}$ hypergeometric and Meijer $G$ functions may be obtained for general $N$ from its definition, but is no more instructive than numerical $z$ moments of eq. (3.5). We show the latter in the right panel of figure -1194 for reference.